40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 1 03:15:03 2014, Jenne, Update, LSC, Yarm locking with CM board Yarm_CMlocking_spectra_30Apr2014_copy.pdfYarm_CMlocking_TFs_30Apr2014_copy.pdf
    Reply  Thu May 1 03:23:07 2014, ericq, Update, LSC, Yarm locking with CM board 
    Reply  Thu May 1 13:03:34 2014, Jenne, Update, LSC, MC board pulled for AC coupling 
       Reply  Thu May 1 14:45:44 2014, Jenne, Update, LSC, MC board back in IMG_1394.JPGIMG_1395.JPGIMG_1398.JPGIMG_1402.JPG
    Reply  Thu May 1 16:41:35 2014, ericq, Update, LSC, Yarm locking with CM board AOsignsSame.pdfAOsignsOpposite.pdfloopModeling.m
       Reply  Thu May 1 17:00:05 2014, rana, Update, LSC, Yarm locking with CM board 
Message ID: 9892     Entry time: Thu May 1 14:45:44 2014     In reply to: 9891
Author: Jenne 
Type: Update 
Category: LSC 
Subject: MC board back in 

Quote:

Quote:

To fix this, I am going to put a 6.8uF cap in series with R30 in the MC board, which is part of the crossbar switch where the IN1 and IN2 get summed.  This should AC-couple the output of the IN2 slider before the summing node.

 MC board is out, so don't be surprised that the MC isn't locking.

 MC board is back in place, MC is locked.

If I disable all of the AO path bits of the CM servo (disable switch, and also gain slider to -32dB), and then move the MC IN2 slider around, the MC does not get an offset anymore (as seen by reduced transmission and increased reflected power), so I think the DC coupling is working.  I do lose lock of the MC if the slider goes above ~22 dB in this situation, but I don't see any effect before then, whereas we were able to see a steady increase in the reflected power as we moved the slider around last night.  So, it seems like things are good with the DC coupling of the IN2 slider.

Here are some photos from before I modified the board (front, back, and zoom of the area I was working in):

IMG_1394.JPG

IMG_1395.JPG

IMG_1398.JPG

And here is my modification, putting the 6.8uF cap in series with (a new) 2k thin film resistor, in the spot for R30:

IMG_1402.JPG

The board is https://dcc.ligo.org/DocDB/0004/D040180/001/D040180-C.pdf

[Edit, 20140721: It looks like this is actually D040180 rev B, not rev C. —Evan]

ELOG V3.1.3-